### v3

# Evaluating Contextualized Representations of (Spanish) Ambiguous Words: A New Lexical Resource and Empirical Analysis

Abstract
Lexical ambiguity—where a single wordform takes on distinct, context-dependent meanings–serves as a useful tool to compare across different language models’ (LMs’) ability to form distinct, contextualized representations of the same stimulus. Few studies have systematically compared LMs’ contextualized word embeddings for languages beyond English. Here, we evaluate semantic representations of Spanish ambiguous nouns in context in a suite of Spanish-language monolingual and multilingual BERT-based models. We develop a novel dataset of minimal-pair sentences evoking the same or different sense for a target ambiguous noun. In a pre-registered study, we collect contextualized human relatedness judgments for each sentence pair. We find that various BERT-based LMs’ contextualized semantic representations capture some variance in human judgments but fall short of the human benchmark. In exploratory work, we find that performance scales with model size. We also identify stereotyped trajectories of target noun disambiguation as a proportion of traversal through a given LM family’s architecture, which we partially replicate in English. We contribute (1) a dataset of controlled, Spanish sentence stimuli with human relatedness norms, and (2) to our evolving understanding of the impact that LM specification (architectures, training protocols) exerts on contextualized embeddings.
Evaluating Contextualized Representations of (Spanish) Ambiguous Words: A New Lexical Resource and Empirical Analysis
Pamela D. Rivière1, Anne L. Beatty-Martínez1, Sean Trott1,2 1Department of Cognitive Science, UC San Diego 2Computational Social Science, UC San Diego {pdrivier, abeattymartinez, sttrott}@ucsd.edu
1 Introduction
State-of-the-art language models (LMs) display a remarkable level of formal linguistic competence (Mahowald et al., 2024). To date, however, we currently lack a precise accounting of the mechanisms underlying LMs’ fundamental linguistic capabilities. The opacity of model internals has motivated work probing the transformations that inputs undergo as they are processed through various model components (Tenney et al., 2019; Hu et al., 2020; Wang et al., 2022; Zou et al., 2023). This work has focused on LMs trained with English-language corpora, with a smaller subset of studies investigating cross-linguistic representations in multilingual models (Chang et al., 2022; Wendler et al., 2024; Michaelov et al., 2023) or comparing representations across multiple monolingual models (Edmiston, 2020). As others have noted Blasi et al. (2022b); Bender (2009), an over-reliance on English as a “model language” limits the generalizability of findings, as well as potential applications Blasi et al. (2022a). Here, we extend interpretability work to Spanish, a language spoken by almost 600M people (with almost 500M native speakers) [^1] https://www.exteriores.gob.es/en/PoliticaExterior/Paginas/ElEspanolEnElMundo.aspx. Specifically, we: (1) evaluate the trajectory of ambiguous Spanish words’ semantic representations within mono- and multilingual LMs and (2) identify layers along a model’s architecture whose semantic representations best capture human judgments of semantic relatedness.
Lexical ambiguity—where a given wordform evokes multiple related or unrelated meanings—offers a unique opportunity to dissociate a word’s form from the contextualized, semantic representations that it can take on as it interacts with a given model’s architecture. Specifically, we can evaluate whether and how LMs integrate the surrounding lexical items in a sentence to produce flexible, context-dependent representations. The representation and processing of ambiguous words is also well-studied in humans Rodd et al. (2004); Martin et al. (1999); Duffy et al. (1988), offering a convenient comparison group. Finally, ambiguity appears to pervade language, with some estimates in English positing that more than 80% of words have multiple meanings (Rodd et al., 2004); more frequent words are also more likely to evoke multiple senses (Zipf, 1945; Piantadosi et al., 2012). Ambiguity, then, is an important phenomenon to contend with in LMs and a useful tool to understand LM representations.
We first provide a survey of related work on ambiguity and interpretability in LMs (Section 2), then present a novel dataset of human relatedness judgments about Spanish ambiguous words—in context (SAW-C) (Section 3). Section 4 documents our use of this dataset to empirically probe the representation of ambiguous words in pre-trained spanish LMs, focusing first on pre-registered analyses [^2] Code and data to run all analyses are available at: https://github.com/seantrott/spanish_norms. Preregistration is available at: https://osf.io/n5htp. of the monolingual BETO (Cañete et al., 2020). We next systematically compare multiple monolingual Spanish and multilingual BERT-based models (Section 5). We conclude with a discussion of the implications of the current results (Section 6), as well as limitations and directions for future work (Section 7).
2 Related work
To date, lexical ambiguity has been largely explored within English monolingual models (Haber and Poesio, 2020; Trott and Bergen, 2021; DeLong et al., 2023). Less work is available with models trained on other languages, such as Spanish (Garcia, 2021; Garí Soler and Apidianaki, 2021), and crucial distinctions emerge across studies within this literature: (1) the operationalization of same and different sense conditions, (2) the degree to which sentential context is controlled around the target word, and (3) the extent to which human semantic judgments are collected to set usage-based expectations for the context-dependence and graded quality of ambiguous word meanings (Erk et al., 2013; Trott and Bergen, 2023).
In Spanish, mono- and multilingual BERT-based models can capture information about semantic relationships between homonyms and their synonyms (Garcia, 2021) and can approximate words’ degree of polysemy (Garí Soler and Apidianaki, 2021). However, available studies tend to leverage naturalistic sentence stimuli from sense-annotated corpora. While valuable, the variability in token sequence length and target word position within naturalistic sentential contexts may make it challenging to isolate the precise effect of the context’s semantic content from the uncontrolled effects of sentence frame (Haber and Poesio, 2020). We thus follow as closely as possible the experimental structure leveraged (for English) in Trott and Bergen (2021), creating Spanish-language sentence pairs that vary along a single context cue evoking either the same or different sense of the target ambiguous noun (see Section 3.1). For our dataset, we document the extent to which context cues presented true minimal pairs (e.g. zero-token-differences across sentence pairs) for the BERT-based models we tested (see Appendix A.1; Table 3).
Using more controlled sentence stimulus design, coupled with empirically collected human benchmarks, prior studies in English have shown that BERT-based models’ contextualized embeddings capture some—though not all—variance in human similarity (Haber and Poesio, 2020) and relatedness (Trott and Bergen, 2021; DeLong et al., 2023) judgments for ambiguous English words. Some of this work has also argued that the continuous nature of LM contextualized representations makes them well-suited as models of human word meanings, which are likely graded to some extent Elman (2009); Li and Armstrong (2024); Li and Joanisse (2021); Trott and Bergen (2023); Rodd et al. (2004); Nair et al. (2020); Zamora-Reina et al. (2022)—though importantly, may also exhibit marked categoriality compared to LM representational spaces Trott and Bergen (2023).
Finally, another important line of work has used techniques like classifier probes Tenney et al. (2019) and activation patching Meng et al. (2022); Wang et al. (2022); Merullo et al. (2024) to decode the putative functional role of different model components (e.g., layers, attention heads, etc.) in producing observed behavior.
To our knowledge, there is little to no work combining these strands of research: i.e., making use of graded human judgments about ambiguous Spanish words to trace the dynamics of contextualized representations in pre-trained Spanish language models. This is the gap we aim to address.
3 Human Annotation Study
We created a dataset containing graded human judgments about Spanish ambiguous words—in context (SAW-C). This process involved first producing and curating materials (i.e., target words and sentences), collecting judgments from native Spanish speakers, and validating those judgments for quality and reliability. This study was pre-registered on the Open Science Framework (OSF) platform.
3.1 Materials
All sentences were developed by two native (Puerto Rican) Spanish speakers, both authors in this study. $102$ target words were drawn from noun lists collected in earlier studies of Spanish lexical ambiguity (Estévez Monzó, 1991; Fraga et al., 2017), as well as spontaneously generated and then verified using the online Real Academia Española [^3] https://www.rae.es/ dictionary. We excluded wordform meanings that corresponded to distinct parts of speech, but we accepted a small fraction ($8/102$) of nouns whose grammatical gender changed across meanings. Each target ambiguous noun was embedded within sentence pairs that differed by a single modifier [^4] For $187/812$ sentence pairs, the modifiers varied along more than just a single word, as was the case when the modifiers required different prepositions, or the grammatical gender of the modifier differed across sentences and required distinct determiners and contractions. Of these, $173$ pairs differed by $1$ word; $14$ differed by $2$ words., termed context cue [^5] Context cues were adjectival modifiers for $100/102$ target nouns; verbs for $2/102$.. The average number of words in sentences was $4.72$. The context cues across the sentence pair could evoke either the Same or Different sense of the word across the contexts.
1a. Compró el aceite de oliva ([S/he] bought the olive oil)
1b. Compró el aceite de cocina ([S/he] bought the cooking oil)
2a. Compró el aceite de motor ([S/he] bought the motor oil)
2b. Compró el aceite de carro ([S/he] bought the car oil)
The minimum number of sentence pairs per target noun was $6$, with a maximum of $28$ (M = $7.96$). A total of $812$ sentence pairs and $102$ target nouns were included in the dataset.
Finally, for the purpose of human norming, the $812$ sentence pairs were assigned to $10$ experimental lists using a Latin Square design, where each list had approximately $81$ or $82$ sentence pairs.
3.2 Participants
Our goal was to collect a minimum of $10$ judgments per sentence pair (i.e., a minimum of $100$ participants). Because we anticipated a non-zero exclusion rate, our pre-registration specified: 1) an initial goal of $120$ participants; and 2) a plan to sample more participants as needed, if any sentence pairs had fewer than $10$ observations after applying the pre-registered exclusion criteria.
Using the two-step process described above, we recruited an initial pool of $139$ participants through Prolific. Participants received \$$\$2.40$ for participating and the median completion time was 12 minutes and 23 seconds, for an average rate of \$$\$11.64$ per hour. On Prolific, we screened for participants who reported that their primary language was Spanish; we specifically recruited participants from the United States, as well as countries in which the dominant language was Spanish (including Chile, México, and Spain).
We excluded participants (1) who failed “catch” trials (where the sentences in the pair were identical), (2) whose task completion times exceeded the sample mean by 3 standard deviations, (3) whose inter-annotator agreement for the items they rated was very low (Pearson’s $r^{2}<0.1$), and (4) who self-identified as non-native Spanish speakers. After all exclusions, we considered data from a total of $131$ participants.
Participants’ self-identified nationalities corresponded heavily to México (69), Spain (39), and Chile (20); 1 participant was from the United States and 2 participants were from Venezuela. [^6] As described in Section 3.4, we found relatively high correlations between mean relatedness judgments produced by participants across Chile, México, and Spain. 54 participants self-identified as female (74 male, 3 non-binary). The average age was $30.97$ (median = $28$), and ranged from $20$ to $59$.
3.3 Procedure
After providing consent, participants were given instructions explaining that some words can have different meanings in different contexts (using an example that was not included in the experimental materials), and that the goal of the experiment was to collect ratings about the relatedness of the meanings expressed by a given word across two sentence contexts.
Figure 1: Sample item from task. Participants emitted graded relatedness judgments from a scale of 1 (totally unrelated) to 5 (same sense) for a given target word (here: aceiteoil), using information provided by the context cue across the sentence pair (here: de oliva / de motorolive / motor).
Each sentence pair was presented on a separate page. Participants were instructed to rate each word on a scale from 1 (totalmente sin relación, “totally unrelated”) to 5 (mismo significado, “same meaning”) [^7] For alternative ordinal annotation schemes in relatedness datasets, see Schlechtweg et al. (2018, 2024, 2025). The target word (e.g., aceite, “oil”) was centered in larger font, and the target sentences were presented side-by-side (the side was randomized across trials). Participants indicated their response via button-press; see Figure 1 for an example. We also included one “catch” trial in the experiment, which simply contained the same sentence, repeated (i.e., the correct answer was $5$).
The entire experiment (including consent form, instructions, and debrief page) was conducted in Spanish. Participants were randomly assigned to lists, and the order of items within each list was randomized. After completing the primary experiments, participants read a debrief form explaining once more that the goal of the experiment was to collect judgments about ambiguous Spanish words, and that their data would be anonymized before analysis and publication.
3.4 Validation of Final Dataset
We validated the final dataset using several approaches. First, we applied multiple exclusion criteria (Section 3.2) and collected a minimum of $10$ ratings per sentence pair. The average number of ratings per pair was $13.1$; the maximum was $17$.
After applying exclusion criteria, we recalculated inter-annotator agreement to estimate the reliability of the ratings in the final dataset. Following past work (Hill et al., 2015; Trott and Bergen, 2021) we calculated inter-annotator agreement using a leave-one-annotator-out scheme. For each of the final $131$ participants, we calculated Spearman’s $\rho$ between the judgments produced by that participant and the mean judgment for those same sentence pairs, leaving out that participant’s data. The resulting distribution of correlation coefficients ranged from $0.39$ to $0.88$, with an average correlation of $0.77$. This number is comparable to past work using similar methods Hill et al. (2015); Trott and Bergen (2021).
Finally, we compared average relatedness judgments across the three main demographic groups reported by participants (Chile, México, and Spain). Judgments across each group were all strongly correlated ($r>0.82$ in all cases).
3.5 Relatedness of Same vs. Different Sense Contexts
We then asked whether and how human relatedness judgments varied as a function of whether two uses of a word (e.g., “aceite”) corresponded to the Same Sense or Different Sense. These will hereafter be considered the levels of the binary variable we call Sense Relationship.
[…]
Figure 2: Density plot representing the distribution of mean relatedness judgments for sentence pairs. As expected, word meanings were rated as more related when used in Same Sense than Different Sense contexts.
4 Analysis of BETO, a Pre-Trained Spanish LLM
Using SAW-C as a probe, we conducted multiple pre-registered analyses on how BETO—a pre-trained monolingual Spanish BERT-based model Cañete et al. (2020)—represents ambiguous words in context, whether and to what extent these representations are predictive of human semantic representations, and which layers of the model contained the most information Tenney et al. (2019). Table 1 summarizes the research questions and their results.
4.1 Model Details
Our pre-registered analyses used the cased version of a Spanish monolingual BERT-based model: BETO, comprised of 12 layers, each made up of 12 self-attention heads, and hidden size of 768, trained on a corpus of approximately 3B words (Cañete et al., 2020) [^8] Accessed via https://huggingface.co/dccuchile/bert-base-spanish-wwm-cased. Exploratory analyses leveraged the models summarized in the Appendix A.1 (Table 2). Each sentence in the dataset (bracketed by special tokens [CLS] and [SEP]) was tokenized separately according to each model’s tokenizer. Sentences contained periods, unlike the sentences viewed by human participants [^9] In a previous version of this study, we presented models with sentences that did not contain periods, to match the stimuli exactly as human participants saw them. However, we found that LMs’ outputs explained more variance in human judgments when the periods were included at the end of the sentences.. We report sentence pair tokenization differences across models in Table 3 (Appendix A.1). When the target noun was represented by multiple subword tokens, we took the average embedding across tokens. We extracted embeddings from each model layer.
4.2 Which layer of BETO best captures sense boundaries?
We first assessed which layers of BETO produced representations that best distinguished between Same Sense and Different Sense uses of word. To address this question, we calculated the Cosine Distance between these contextualized representations of the target word from each sentence pair for each layer. Concretely, this yielded $812$ Cosine Distance values for each layer of BETO.
Then, we asked how Cosine Distance evolved through the network’s layers with respect to the Same/Different Sense distinction. We built a series of logistic regression models in R with Sense Relationship as a dependent variable, and Cosine Distance from a given layer $\ell_{i}$ of BETO as an independent variable. We then measured the Akaike Information Criterion (or $AIC$) Akaike (2011); Burnham and Anderson (2004) of the resulting model as a measure of model fit. The best-fitting model used Cosine Distance from layer 6. Figure 3 highlights the change in Cosine Distance across layers of BETO as a function of Sense Relationship, suggesting that the difference between conditions was largest at this layer.
Figure 3: Average Cosine Distance between the contextualized representations of the target ambiguous word across each layer of BETO, depicted as a function of whether the contexts cued the Same Sense or Different Sense.
4.3 Which layer of BETO best predicts relatedness?
Our second question was whether certain layers of BETO produced representations that better predicted human relatedness judgments than others. For each layer, we calculated the correlation coefficient (both Pearson’s r and Spearman’s $\rho$) between Cosine Distance values obtained from BETO and the distribution of Mean Relatedness judgments obtained for each sentence pair. We also calculated $R^{2}$ as an estimate of the amount of variance explained in human relatedness judgments as a function of Cosine Distance from that layer alone.
[…]
4.4 BETO under-performs inter-annotator agreement
We then compared the best-performing layer ($\ell=12$) to human inter-annotator agreement (calculated in Section 3.4). The correlation magnitude of the best-performing layer ($\rho=0.61$) was considerably lower than the average inter-annotator agreement ($\bar{X}_{\rho}=0.77$), with BETO’s performance lying approximately in the bottom $5\%$ of the distribution of agreement values (Figure 4).
Figure 4: Distribution of human inter-annotator agreement scores, calculated using a leave-one-annotator out scheme. The vertical dashed line represents the correlation between human judgments and Cosine Distance values extracted from BETO.
4.5 BETO is Less Sensitive to Sense Boundaries
Past work conducted in English Trott and Bergen (2021, 2023) suggests that LMs are less sensitive to sense boundaries—the distinction between Same and Different Sense—than humans. However, it is unclear whether this effect generalizes to Spanish speakers and Spanish LMs.
[…]
Finally, we extracted the residuals of a linear regression model predicting Mean Relatedness from Cosine Distance alone. We then plotted the distribution of these residuals according to Sense Relationship. As illustrated in Figure 5, BETO (as well as all the Spanish language models tested in Section 5) consistently underestimated the relatedness of Same sense pairs, and consistently overestimated the relatedness of Different sense pairs.
Figure 5: Residuals of linear regression models fit for each LM, predicting relatedness from the interaction between cosine distance and layer position; residual distributions are separable as a function of Sense Relationship.
Research Question Result Section
Do humans judge Same Sense uses to be more related than Different Sense uses? Yes 3.5
Which layer of BETO is most sensitive to the Same/Different Sense distinction? $\ell=6$ 4.2
Which layer of BETO is most correlated with human relatedness judgments? $\ell=12,7$ 4.3
Does BETO match human inter-annotator agreement? No 4.4
Does BETO “explain away” the effect of categorical Sense Relationship in humans (e.g. “sense boundaries”)? No 4.5
Table 1: Summary of pre-registered research questions and their results.
5 Comparing Pre-Trained Spanish Language Models
Section 4 tested several pre-registered hypotheses (Table 1) with respect to a single pre-trained Spanish LM. Here, we extend this work in exploratory analyses of additional pre-trained Spanish BERT-based models. Testing multiple models is an important step towards establishing the external validity of a finding; additionally, it is useful for testing hypotheses about model scale Kaplan et al. (2020) or other model specifications (e.g., architecture, multilingual status).
5.1 Models
We considered $10$ monolingual Spanish language models (including BETO) and $2$ multilingual models. Models varied in their training procedures (e.g., BERT vs. RoBERTa; Liu et al. (2019)), tokenization scheme, number of layers, training corpus size, total number of parameters, and whether or not they were multilingual (Table 2).
5.2 Impact of model scale
Past work Kaplan et al. (2020) suggests that increases in a model’s number of parameters may correlate with metrics of model performance (e.g., perplexity). At the same time, there is some evidence that increasing scale does not always produce more human-like representations, i.e., large models with lower perplexity do not always better predict human reading times Kuribayashi et al. (2021).
To assess this question, we compared the maximum $R^{2}$ achieved by each of the $12$ language models [^10] I.e., the $R^{2}$ from the best-performing layer of each model. and asked whether a model’s best $R^{2}$ was related to its size. We found that model size was in fact correlated with its ability to predict human relatedness judgments in Spanish  [^11] Though, we note that this becomes true only when models are tested with stimuli containing periods at the end of the sentences. The RoBERTa model family was particularly sensitive to this implementational difference. (Figure 6). RoBERTa-large (with the most parameters) outperformed all other models. Multilingual models, however, performed just as poorly as monolingual models with many fewer parameters. Notably, models within the Spanish-language ALBERT family—carefully controlled for various training and architectural properties Cañete et al. (2022)—failed to show clear, consistent evidence of scaling beyond the jump from -tiny to -base architectures. None of the models matched the variance explained in average judgments by individual human annotators (i.e., inter-annotator $R^{2}$).
Figure 6: Maximum $R^{2}$ achieved by each model by number of parameters and multilingual status. Horizontal dashed line depicts the average variance explained in a leave-one-annotator-out scheme.
5.3 Performance across layers
In Section 4, we found that cosine distance measures extracted from the middle and last layers of BETO were most useful for predicting whether two contexts belonged to the same sense, and for predicting human relatedness judgments. Do other models or model families show the same trajectory of performance across layers?
Figure 7: Depiction of seven pre-trained Spanish LMs’ ability to predict human relatedness judgments across layers (measured as $R^{2}$). For ease of illustration, this plot shows only LMs with $12$ layers.
Models varied in their number of layers. Thus, we first compared the trajectory of $R^{2}$ across layers for the subset of models with the same number of layers, i.e., $12$ layers (Figure 7) or $24$ layers (Figure 9). In each case, we identified two qualitatively distinct “classes” of trajectory: a rise and plateau trajectory, in which performance improves up until a point (e.g., layer 6) and then stays relatively stable; and a rise and fall trajectory, in which performance improves and then decays substantively in the final layers (Figure 8).
In order to compare all models on the same axis, we calculated the layer depth ratio, which divides each layer position in a given network by the total number of layers in that network. We then visualized the average $R^{2}$ by layer depth ratio across the three model families tested: ALBERT, BERT, and RoBERTa. As Figure 8 suggests, the two putative trajectories appear to covary with model family: the ALBERT family of models shows a rise and fall trajectory, while the BERT and RoBERTa family of models shows a rise and plateau trajectory.
Figure 8: Depiction of LM ability to predict mean relatedness judgments (measured as $R^{2}$), broken down by Model Family and Layer Depth Ratio, i.e., with each layer divided by the the total number of layers in a given model. 4 BERT models (1 multilingual), 3 RoBERTa models (1 multilingual), and 5 ALBERT models.
5.4 Sensitivity to sense categoriality across models
Lastly, we asked whether the LMs tested explained away the sense boundary effect in humans. Each model showed the same pattern as BETO: Cosine Distance underestimated the relatedness of Same Sense meanings and overestimated the relatedness of Different Sense meanings (Figure 5).
6 Discussion
We introduced a novel dataset (SAW-C) containing human relatedness judgments about ambiguous Spanish words in controlled, minimal pair contexts. Using this dataset, we probed pre-trained BERT-based models’ representations of ambiguous words, finding that: 1) LM representations correlate with human judgments but do not match inter-annotator agreement, and exhibit systematic errors; 2) performance varies across layers, with model families showing distinct trajectories of performance; and 3) performance scaled with model size.
The systematic underestimation and overestimation errors observed with respect to same vs. different sense contexts (Figure 5) is consistent with past work conducted in English (Trott and Bergen, 2021, 2023). One explanation for this is that the initial (static) embedding for an ambiguous wordform might entangle all of its multiple meanings Grindrod (2024), which must then be “teased apart” in context—but which might nevertheless persist as “attractors” in subsequent layers. Disambiguation could be made even more difficult by the presence of minimal pair contexts Garcia (2021).
Psycholinguistic research suggests that humans also activate uncued, dominant meanings in certain tasks Duffy et al. (1988); Martin et al. (1999); here, however, humans appeared to distinguish target meanings with relative ease. It is possible that humans represent distinct homonymous (though not necessarily polysemous) meanings along clearly differentiable regions of meaning space (Rodd et al., 2004; Trott and Bergen, 2023, 2021; Haber and Poesio, 2020), which would be consistent with the fact that categorical sense boundaries (Same vs. Different sense conditions) explained an overwhelming share of human relatedness judgments (Section 4.5, Trott and Bergen (2023)).
Lastly, this work contributes to expanding the linguistic diversity of both human-annotated benchmarks and interpretability research. Although Spanish ranks among the most widely spoken languages, it suffers from a surprising dearth of resources (with notable exceptions, Baldissin et al. (2022)), pre-trained models, and interpretability research—particularly when compared to English. In one study, a sample of 550 corpora (spanning 22 languages) contained >$50\%$ English-language corpora, while <$10\%$ represented Spanish-language corpora Anand et al. (2020). A wider research perspective—considering varied languages—is critical for ensuring the generalizability of findings.
7 Limitations
7.1 Limitations of the Dataset
SAW-C is limited in scope, containing only $812$ sentence pairs. This is considerably smaller than English benchmarks such as BLiMP Warstadt et al. (2020), which contain tens of thousands of examples. However, it is larger than or comparable in size to other, more targeted datasets involving crowd-sourced human annotations Erk et al. (2013); Haber and Poesio (2020); Trott and Bergen (2021). Relative to specifically Spanish-language datasets, ours is the only one we are aware of that collects human judgments for target ambiguous words embedded within minimal pair stimuli [^12] For examples of other Spanish-language datasets that collect human judgments under different experimental conditions, see Estévez Monzó (1991); Gómez-Veiga et al. (2010); Domínguez et al. (2001); Haro et al. (2017); Fraga et al. (2017); Zamora-Reina et al. (2022); Baldissin et al. (2022).. Importantly, SAW-C includes not only the sentence pairs but also over $10000$ validated human judgments from $131$ participants about those sentence pairs.
Another limitation concerns stimulus generation. The ambiguous nouns included in our dataset were spontaneously produced by native (Caribbean) Spanish speakers, or selected from previously published lists (Estévez Monzó, 1991; Fraga et al., 2017), rather than via automated searches. Spontaneous production of ambiguous words may (1) overrepresent homonymous words (Estévez Monzó, 1991), and (2) underrepresent words whose multiple meanings have large dominance asymmetries Duffy et al. (1988). In future work, we intend to collect dominance norms for these items. Differences in Spanish varieties may have also led some items to appear awkward to annotators from different nationalities (Baldissin et al., 2022; Lipski, 2014). We address this concern by showing strong correlations in relatedness judgments from the three heavily represented groups (Chile, Spain, and México, Section 3.4), but other groups’ relatedness judgments may not be represented in this sample.
Finally, our sentence pairs are not naturalistic. Our controlled minimal pair design was intentional, enabling us to identify key differences between LM and human representations, e.g., LMs display less sensitivity than humans to the manipulation of word meaning across minimal pair contexts (Figure 5). At the same time, it is important to know whether and to what extent the current results replicate with naturalistic stimuli. Thus, we aim to augment SAW-C with naturalistic examples of the target ambiguous words.
7.2 Limitations of the Analysis
All the analyses presented here are essentially correlational. As others have noted Hewitt and Liang (2019); Niu et al. (2022); Zhou and Srikumar (2021), supervised methods for probing LM representations are more informative about the ability of the probe to learn specific features than the question of whether the LM “naturally” encodes that feature and deploys it for token prediction. Future work would benefit from the selective application of “knock-out” methods or “activation patching” Meng et al. (2022), both of which have proven more successful in characterizing the causal, mechanistic role of model components. We view the current work as a useful starting point, which can motivate future work isolating the mechanistic role of specific model circuits within each layer.
The finding that distinct model families exhibit distinct trajectories in performance ( Figure 8) is intriguing, but due to its exploratory nature, it is unclear to what extent this finding is reliable and robust to different datasets or probing methods. In a supplementary analysis (see Appendix 11), we found qualitatively similar clusters of trajectories in pre-trained English models, though these differences appeared considerably weaker than in the Spanish models. Future work could build on the question of whether—and more importantly, why— distinct model architectures and training schemes lead to different processing mechanisms.
We also note that analysis of performance as a function of model size was not an ideal test of the scaling hypothesis, given that many features of these models (e.g., training data, model architecture) were not controlled, except within the Spanish ALBERT family.
Lastly, the structure of Spanish modifiers (which follow the target noun) meant that we could only test encoder (e.g. bidirectional) models in this study. To offset this limitation, we tested a large suite of Spanish-trained encoder models, but future studies may leverage stimuli modified to suit decoder-only architectures.
8 Ethical Considerations
This research was conducted with IRB approval. All data from human participants has been fully anonymized before analysis and publication.
Acknowledgments
Pamela D. Rivière is supported by UCSD’s Chancellor’s Postdoctoral Fellowship Program. We thank Benjamin Bergen and Cameron Jones for valuable discussions.
Appendix A Appendix
A.1 LLM Specifications and Dataset Tokenization
See Table 2 for a summary of the LMs considered in Section 5, including architecture, multilingual status, corpus size, tokenization scheme, training objective, number of layers, and number of parameters. Because models used different tokenizers, we also calculated summary statistics about the number of tokens in each sentence in each sentence pair, as well as the average number of token differences (i.e., 5 vs. 4 tokens across the members of a given pair) for each tokenizer (see Table 3).
Model # Lang Corpus Tok. Trn Obj # Layers # Params
BETO 1 $\sim$ 3B SP DM, WWM 12 $\sim$ 110M
BETO-uncased 1 $\sim$ 3B SP DM, WWM 12 $\sim$ 110M
mBERT 104 ? WP MLM, WWM?, NSP 12 $\sim$ 178M
DistilBETO 1 $\sim$ 3B SP DistilLoss, MLM 6 $\sim$ 66M
ALBETO-tiny 1 $\sim$ 3B SP MLM 4 $\sim$ 5M
ALBETO-base 1 $\sim$ 3B SP MLM 12 $\sim$ 12M
ALBETO-large 1 $\sim$ 3B SP MLM 24 $\sim$ 18M
ALBETO-xlarge 1 $\sim$ 3B SP MLM 24 $\sim$ 59M
ALBETO-xxlarge 1 $\sim$ 3B SP MLM 12 $\sim$ 223M
RoBERTa-BNE-base 1 $\sim$ 135B byte-BPE MLM (DM?) 12 $\sim$ 125M
RoBERTa-BNE-large 1 $\sim$ 135B byte-BPE MLM (DM?) 24 $\sim$ 355M
XLM-RoBERTa 100 ? SP MLM 12 $\sim$ 278M
Table 2: Spanish language model properties and training procedures. Models are cased (distinguish between upper and lowercase characters) unless otherwise specified. All monolingual models are trained on Spanish-language corpora; multilingual models include Spanish-language corpora. Model Notes: For mBERT, (a) the corpus size per language varied and we are unsure of the total corpus size13, (b) it was unclear to us whether the current version of the model on HuggingFace is updated with the whole-word masking (WWM) technique14 during pre-training. For RoBERTa-BNE models, it was unclear to us whether authors used dynamic masked (DM) modeling, as in the English RoBERTa. For XLM-RoBERTa, the corpus size per language may have varied, but we were uncertain to what extent pretraining text was sampled proportionally to its representation in the corpus (Conneau et al., 2020). Acronyms: BNE: Biblioteca Nacional de España (National Library of Spain); byte-BPE: byte-level Byte-Pair Encoding; DistilLoss: Distillation loss (Sahn et al., 2019; Cañete et al., 2022); DM: Dynamic Masking (Liu et al., 2020); MLM: Masked Language Modeling; NSP: Next Sentence Prediction; SP: SentencePiece; WWM: Whole-Word Masking; WP: WordPiece. This summary represents our best attempt at gathering and reconstructing some model specifications—they are necessarily incomplete, and may contain inaccuracies borne from either a lack of knowledge regarding more recent updates or an imperfect understanding of the training protocols as described in the relevant primary literature and repositories.
1313footnotetext: See https://github.com/google-research/bert/blob/master/multilingual.md for available details on language-specific corpus size selection.1414footnotetext: See https://github.com/google-research/bert for notes on the English BERT pre-training update using whole-word masking.
Model Avg ; Modal ; Max Target Noun
Token Differences # Tokens
BETO $\sim$ 0.682 ; 1 ; 3 $\sim$ 1.07
BETO-uncased $\sim$ 0.670 ; 1 ; 3 $\sim$ 1.03
mBERT $\sim$ 0.997 ; 1 ; 4 $\sim$ 1.27
DistilBETO $\sim$ 0.670 ; 1 ; 3 $\sim$ 1.03
ALBETO-tiny $\sim$ 0.619 ; 1 ; 3 $\sim$ 1.04
ALBETO-base $\sim$ 0.619 ; 1 ; 3 $\sim$ 1.04
ALBETO-large $\sim$ 0.619 ; 1 ; 3 $\sim$ 1.04
ALBETO-xlarge $\sim$ 0.619 ; 1 ; 3 $\sim$ 1.04
ALBETO-xxlarge $\sim$ 0.619 ; 1 ; 3 $\sim$ 1.04
RoBERTa-BNE-base $\sim$ 0.643 ; 0 ; 3 $\sim$ 1.05
RoBERTa-BNE-large $\sim$ 0.643 ; 0 ; 3 $\sim$ 1.05
XLM-RoBERTa $\sim$ 0.929 ; 1 ; 4 $\sim$ 1.28
Table 3: Average, modal, and maximum token differences across sentence pairs per LM. Tokenization schemes for all Spanish monolingual LMs were heavily represented by either non-zero or single-token differences across sentence pair stimuli, whereas the multilingual models tested here tended to more frequently generate non-zero token differences across the sentence pairs.
A.2 Analysis of Expected Layer
In the primary manuscript, we identified which layers of BETO provided the best fit, i.e., which were most effective for predicting Sense Relationship (layer 6) and which were most effective at predicting Mean Relatedness judgments (layer 12). However, in some cases, the improvements across layers are fairly marginal. Thus, in this supplementary analysis, we implemented a version of the Expected Layer analysis described by Tenney et al. (2019). This analysis considers the size of the improvement across layers and estimates the layer at which particular kinds of information is expected to resolve in the network.
A.2.1 Methods
The Expected Layer statistic considers the improvement in performance (measured here as $AIC$ or $R^{2}$, depending on the analysis in question) across progressively more complex regression models fit with cosine distance information from each layer. This improvement in performance measure was defined as:
$\centering\Delta^{\ell}=Score(P_{T}^{\ell})-Score(P_{T}^{\ell-1})\@add@centering$ (1)
Where $Score(P_{T}^{\ell})$ is defined as the performance ($AIC$ or $R^{2}$) of a regression model equipped with cosine distance information from a given layer $\ell$ and each previous layer, i.e., such that the number of parameters in the regression model was equal to $\ell$. (Note that this was distinct from the approach taken in the primary manuscript, in which distinct univariate regression models were fit for each layer.) The Expected Layer statistic itself was defined as follows:
$\centering\bar{E}_{\Delta}[\ell]=\frac{\sum_{\ell=1}^{L}\ell\cdot\Delta^{(\ell)}_{T}}{\sum_{\ell=1}^{L}\Delta^{(\ell)}_{T}}\@add@centering$ (2)
A.2.2 Results
Using this approach, we obtained Expected Layer statistics for both predicting Sense Relationship ($3.6$) and for predicting Mean Relatedness ($3.4$). Note that in both cases, the Expected Layer was smaller than the layer at which optimal performance was achieved; this is consistent with the observation that past a certain point, additional LM layers resulted in only marginal gains in prediction.
Figure 9: Depiction of three pre-trained Spanish LMs’ ability to predict human relatedness judgments across layers (measured as $R^{2}$). For ease of illustration, this plot shows only LMs with $24$ layers.
A.3 Additional analyses of pre-trained English BERT-based models
In the primary manuscript, we reported the results of work using ambiguity as a probe for understanding and interpreting how pre-trained Spanish LMs process word meanings. We found that larger models exhibited better performance; and secondly, that different model families exhibited different trajectories of performance across layers. We asked whether these results replicated in pre-trained English models, using an openly available dataset of relatedness judgments about ambiguous English words, in context (RAW-C) Trott and Bergen (2021).
A.3.1 A correlation between model scale and performance
Figure 10: Pre-trained monolingual English language models show evidence of scaling, i.e., models with more parameters achieve a higher $R^{2}$ in predicting human relatedness judgments about ambiguous English words.
[…]
A.3.2 Layer-wise trajectories by model family
We also asked whether different model families displayed different performance trajectories across layers. In the primary manuscript, we found that the ALBERT family models displayed a rise and fall trajectory, while both BERT and RoBERTa displayed rise and plataeau trajectories (see Figure 8). Surprisingly, we found qualitatively similar (albeit weaker) classes of trajectories using the pre-trained English models on the RAW-C dataset (see Figure 11).
Figure 11: Depiction of English LMs’ ability to predict mean relatedness judgments (measured as $R^{2}$), broken down by Model Family and Layer Depth Ratio, i.e., with each layer divided by the the total number of layers in a given model.
A.4 Analysis of GPT-4 Turbo
In the primary manuscript, we found that the Large Language Models tested produced representations that were correlated with human judgments, but nonetheless systematically underestimated how related people judged Same Sense meanings to be—and overestimated how related people judged Different Sense meanings to be (see Figure 5). However, recent work Trott (2024); Dillion et al. (2023) suggests that state-of-the-art LLMs like GPT-4 are capable of producing “norms” that accurately predict human judgments across various domains, including the relatedness of English words.
Because these models are “closed source”, this work typically relies on prompting the models with instructions and directly eliciting a judgment (e.g., a relatedness rating). Thus, a key limitation is that even if these judgments are highly correlated with human judgments, it is very difficult (in some cases impossible) to know why, i.e., which representations or mechanisms give rise to the behavior in question—making them less well-suited to questions about model interpretability.
Nonetheless, the question of the empirical fit between these LM judgments and human judgments is still an interesting one—particularly because past work has primarily focused on judgments in English, and it is unclear whether these LMs would excel at other languages. Thus, in this supplementary analysis, we asked whether GPT-4 Turbo, a state-of-the-art LM, produced judgments that were more predictive of human judgments than the models tested in the primary manuscript.
A.4.1 Methods
Following past work Martínez et al. (2024); Trott (2024); Dillion et al. (2023), we prompted GPT-4 Turbo (gpt-4-1106-preview) using the OpenAI Python API. GPT-4 Turbo was presented with a system prompt containing the same instructions (in Spanish) that were presented to human participants, explaining the purpose of the task. Then, for each sentence pair, Turbo was presented with the same instructions given to human participants (again in Spanish) asking them to rate the relatedness of the target word across the two contexts. The two sentences were presented on separate lines, as was the target word (e.g., “Word: aceite”). Finally, we included an additional instruction requesting a single number in response. Turbo was prompted using a temperature of 0 and its responses were limited to a maximum of 3 tokens.
A.4.2 Results
The ratings produced by Turbo were highly correlated with human judgments, approaching or even exceeding average human inter-annotator agreement ($\rho=0.79$).
[…]
As in Figure 5, we also visualized the residuals of a linear model containing only Rating as a predictor (Figure 12). Notably, even though Rating was highly correlated with Mean Relatedness, the residuals suggest that Turbo’s ratings follow a similar pattern with respect to sense boundaries as was observed with BETO and the other models tested: GPT-4 consistently underestimates the relatedness of Same Sense pairs, and consistently overestimates the relatedness of Different Sense pairs.
Figure 12: Residuals from a linear regression predicting Mean Relatedness from ratings elicited from GPT-4 Turbo. Although Turbo’s ratings are highly correlated with human judgments, they still systematically underestimate the relatedness of Same Sense pairs and overestimate the relatedness of Different Sense pairs.
A.4.3 Discussion
In this supplemental analysis, we found that a larger model trained with Reinforcement Learning from Human Feedback (RLHF) produced behavior that was much more correlated with human relatedness judgments than the other Spanish language models tested.
On the one hand, this analysis has considerable limitations: as noted above, a major challenge with relying on large language models such as GPT-4 is that, despite their impressive performance, much is still unclear about how exactly they were trained—either in terms of the original training data or the procedure for implementing RLHF. Thus, as a tool for testing scientific hypotheses, they may be less useful than open-source models, i.e., the ones tested in the primary manuscript, especially if the scientific question under investigation concerns the representations or mechanisms used by the LLM. These are, of course, the kinds of questions that research on interpretability is generally interested in.
On the other hand, the fact that an LLM produces behavior that rivals inter-annotator agreement suggests that the representations required to produce this behavior can be learned provided sufficient training data and fine-tuning; in this case, direct data contamination is an unlikely concern given that the materials were entirely novel and the ratings had never been published before. This analysis is also notable in that it reveals that even a very large-scale model trained with RLHF appears to be less sensitive to sense boundaries than human judgments, as depicted in Figure 12.
However, because Turbo is a closed-source model, it is difficult to draw firm conclusions from these results precisely because we cannot investigate where or how this behavior emerges. For example, we cannot investigate which layer of GPT-4 Turbo appears to be most helpful for producing high-quality relatedness judgments. Future work would thus benefit from investigating which architectural features, training objectives, or fine-tuning procedures are likely candidates for producing this improvement in performance, ideally in open-source Spanish language models.
A.5 Additional sentence stimuli examples
All sentence stimuli will be made available following publication. Here, we list additional examples from the dataset. Since target ambiguous nouns in Spanish are not always ambiguous in English, we mark the target noun’s English translation in bolded italics within each parenthetical.
1a. Era un banco financiero. (It was a financial bank.)
1b. Era un banco exitoso. (It was a successful bank.)
2a. Era un banco cómodo. (It was a comfortable bench.)
2b. Era un banco de plástico. (It was a plastic bench.)
1a. Disfrutó el baño caliente. ([S/he] enjoyed the warm bath.)
1b. Disfrutó el baño frío. ([S/he] enjoyed the cold bath.)
2a. Disfrutó el baño remodelado. ([S/he] enjoyed the remodeled bathroom.)
2b. Disfrutó el baño privado. ([S/he] enjoyed the private bathroom.)
1a. Tenía una muñeca fracturada. ([S/he] had a fractured wrist.)
1b. Tenía una muñeca lesionada. ([S/he] had an injured wrist .)
2a. Tenía una muñeca preciosa. ([S/he] had a lovely doll.)
2b. Tenía una muñeca plástica. ([S/he] had a plastic doll.)